# Koopman Autoencoders with Continuous-Time Latent Dynamics for Fluid Dynamics Forecasting

### v1

@@ Abstract
Data-driven surrogate models have emerged as powerful tools for accelerating the simulation of turbulent flows. However, classical approaches which perform autoregressive rollouts often trade off between strong short-term accuracy and long-horizon stability. Koopman autoencoders, inspired by Koopman operator theory, provide a physics-based alternative by mapping nonlinear dynamics into a latent space where linear evolution is conducted. In practice, most existing formulations operate in a discrete-time setting, limiting temporal flexibility. In this work, we introduce a continuous-time Koopman framework that models latent evolution through numerical integration schemes. By allowing variable timesteps at inference, the method demonstrates robustness to temporal resolution and generalizes beyond training regimes. In addition, the learned dynamics closely adhere to the analytical matrix exponential solution, enabling efficient long-horizon forecasting. We evaluate the approach on classical CFD benchmarks and report accuracy, stability, and extrapolation properties.
@@ 1 Introduction
Deep learning has recently emerged as a promising tool for flow prediction (Borrelli et al., 2022; Srinivasan et al., 2019; Fukami et al., 2020; Morimoto et al., 2021). Classical data-driven solvers operate directly on the data space, mapping flow fields at the current timestep to the next autoregressively using deep learning models such as Convolutional Neural Networks (CNNs) (Krizhevsky et al., 2012), Graph Neural Networks (GNNs) (Kipf and Welling, 2017) and Neural Operators (Lu et al., 2021; Li et al., 2021). While such data-space models excel at capturing fine-scale features, they tend to suffer from error accumulation, high computational cost due to step-by-step rollouts, and reduced stability over long horizons.
Alternatively, latent-space models compress high-dimensional flow fields into low-dimensional embeddings and evolve them over time using different network structures. LE-PDE, proposed by Wu et al. (2022), illustrates that learning the evolution of dynamics in a global latent space is a simple, fast and scalable method to accelerate the simulation. Within the latent space paradigm, a class of physics-based methods based on Koopman theory (Koopman, 1931) has been proposed. Koopman theory states that the nonlinear dynamics of a system could be represented as an infinite-dimensional linear operator, known as the Koopman operator. By linearizing the dynamics, Koopman theory enables the use of spectral analysis and linear algebra to study nonlinear dynamics. However, identifying the Koopman operator is challenging due to its infinite-dimensional nature. One classical approach is Dynamic Mode Decomposition (DMD) (Kutz et al., 2016), which estimates finite-dimensional approximations of the Koopman operator directly from data. To improve its expressive power, Takeishi et al. (2017) extended DMD by introducing neural networks to learn nonlinear feature maps, enabling Koopman operator estimation in a learned feature space. Building on this line of work, Lusch et al. (2018) proposed an encoder-decoder architecture, forming the basis of Koopman Autoencoders (KAEs).
Latent-space evolution reduces computational cost and improves long-term stability, but most existing approaches rely on discrete-time updates, which can limit flexibility and temporal generalization. While effective, this formulation constrains the model to the sampling resolution of training data and can hinder extrapolation across different timesteps. Inspired by recent advances in neural ODEs, we propose to use a continuous-time update for latent space evolution. By modelling the latent dynamics through an ODE and integrating them with stable numerical schemes, our approach decouples the learned dynamics from fixed time intervals, enabling flexible rollouts, improved robustness, and principled generalization across temporal scales. We compare our proposed methodology with a benchmark work using conditional diffusion models (Kohl et al., 2026) in terms of accuracy and efficiency. Our contributions are summarized below.
@@
  • Continuous-Time Koopman Autoencoder: We introduce a Koopman autoencoder framework where latent dynamics are modeled in continuous time, enabling stable and flexible temporal evolution beyond the training timestep.
  • Physics-Informed, Parameter-Dependent Latent Dynamics: The latent Koopman operator is conditioned on physical parameters (e.g., Reynolds or Mach numbers), allowing the model to generalize across different flow regimes while maintaining physically consistent predictions.
  • Latent Dynamics Consistency: We demonstrate that the learned latent dynamics closely match the analytical solution via matrix exponentiation, enabling efficient and accurate long-horizon forecasting without iterative rollouts.
  • Efficiency and Robustness: Our approach significantly reduces computational cost compared to autoregressive diffusion models, while supporting zero-shot generalization across variable timesteps.
@@ 2 Related Work
Autoregressive prediction in data space. Classical deep learning approaches for flow forecasting adopt autoregressive prediction in the data space. CNNs, particularly U-Net architectures (Ronneberger et al., 2015), have been widely used to capture spatial features in an image-to-image translation framework (Deo et al., 2023; Wang et al., 2019). Other approaches have also explored generative models. For instance, Generative Adversarial Networks (GANs) (Goodfellow et al., 2020) have been employed to predict complex unsteady flow dynamics around objects like cylinders (Lee and You, 2019). Similarly, conditional diffusion models (Batzolis et al., 2021) have been benchmarked for their ability to predict turbulent flow simulations autoregressively (Kohl et al., 2026).
Latent-space autoregressive models. Many works compress flow fields into a latent representation and evolve it autoregressively. This approach decouples the spatial complexity from the temporal dynamics. Classical examples include CNN autoencoders coupled with LSTMs (Hochreiter and Schmidhuber, 1997) and Recurrent Neural Networks (RNNs) to propagate the latent state forward in time (Eivazi et al., 2020; Hasegawa et al., 2020; Bukka et al., 2020; Maulik et al., 2021; Zhou and Cheng, 2025). Similarly, Graph Neural Networks (GNNs) have proven effective for systems with irregular geometries to conduct temporal evolution in latent space (Sanchez-Gonzalez et al., 2020). More recently, transformers (Vaswani et al., 2017) have also been adapted for this task, leveraging its powerful attention mechanism to capture long-range temporal dependencies in the latent dynamics of fluid flows (Hemmasian and Farimani, 2023).
Koopman autoencoders. Building upon (Lusch et al., 2018)’s foundational work, subsequent studies have introduced enhancements. Azencot et al. (2020) proposed Consistent Koopman Autoencoders, integrating both forward and backward dynamics and enhancing accuracy for long time predictions. Temporally-Consistent Koopman Autoencoder (tcKAE) (Nayak et al., 2025) achieved accurate long-term predictions with limited and noisy training data through a consistency regularization term. Halder et al. (2026) proposed Koopman $\beta$-variational autoencoders for reduced-order modeling (ROM) of turbulent flows. $K^{2}$VAE integrates KalmanNet with KAEs to refine predictions and model uncertainty (Wu et al., 2025).
@@ 3.1 Koopman Autoencoders
Koopman operator theory provides a linear perspective on nonlinear dynamics. Consider a discrete-time dynamical system with states $x_{t}\in\mathbb{R}^{N_{d}}$, where $x_{t}$ denotes the state of the system at time $t\in N$. The state evolves to its next timestep $t+\Delta t$ according to a flow map $f:\mathbb{R}^{N_{d}}\to\mathbb{R}^{N_{d}}$:
@@
which approximates the set of observables $h(x_{t})$. In this latent space, the evolution can be modelled as an approximately linear finite-dimensional Koopman operator $\mathbf{K}$,
@@
which naturally accommodates variable timesteps, enabling variable-step integration and extrapolation beyond training timescales.
3.2 Numerical Integration and Latent Dynamics
In contrast to Neural ODEs which model dynamics using black-box models, our approach restricts latent dynamics to a linear form, thereby aligning with the linear latent dynamics described by Koopman theory. This linearity also enables precise state evolution through matrix exponentiation, yielding efficient long-horizon forecasting.
Linear Stability and Integration.
In our method, we need to choose a numerical integration scheme for state evolution. Forward Euler integrators are simple to implement and interpret, however, they tend to diverge unless the integrator step is small enough, thus making them unpractical. Among available choices, the classical fourth-order Runge–Kutta (RK4) method offers a strong balance of accuracy, stability and interpretability, making it particularly suitable for long-rollout predictions.
Integration via Matrix Exponentiation.
During inference time, we can leverage the analytical solution of the learned linear system to bypass iterative solvers entirely. Given the linear system in (6), the exact solution for any future state $z_{\tau}$ given $z_{0}$ is defined by the matrix exponential $z_{\tau}=\exp(\mathbf{K}\tau)z_{0}$, where we can predict the state at arbitrary time $\tau$ directly, independent of the step size $\Delta t$. The matrix exponential $\exp(\mathbf{M})$ is defined by power series:
@@ 4.1 Proposed Model
Let $x_{t}\in\mathbb{R}^{N_{d}}$ denote the high-dimensional physical state of the system. We assume the data is at discrete timestamps $t_{i}$, which are uniformly spaced by a fixed time interval $\Delta t$, such that $t_{i+1}=t_{i}+\Delta t$. Accordingly, $x_{t_{i}}$ represents the state snapshot at time $t_{i}$.
@@ 4.1.1 Encoder and Decoder
We employ a Transformer-based encoding architecture which is split into two streams: a history encoder ($\mathcal{E}_{\text{history}}$) and a present encoder ($\mathcal{E}_{\text{present}}$). The history encoder is designed to process a sequence of past states. While the architecture supports an arbitrary number of previous steps, for consistency with benchmark evaluations, we employ an Auto-Regressive order of 2 ($AR\text{-}2$). Suppose the current state is $x_{t_{i}}$, the history encoder conditions only on the immediate past state $x_{t_{i-1}}$. The present encoder processes the current state $x_{t_{i}}$. The initial latent state $z_{t_{i}}$ is constructed by taking the average of outputs from two encoders,
@@ 4.1.2 Koopman Operator
Next, we discuss the structure of the Koopman operator. A key innovation in our approach is the dependence of the Koopman operator on external control parameters, denoted by $\phi$. In physical systems, such as turbulent flows, $\phi$ may represent the Reynolds number or other governing coefficients. Unlike traditional Koopman approaches where the operator is constant across the domain, our operator varies based on the physical regime.
@@
where RK4 integration is used.
Finally, we define the structure of the operator. To capture the dependency on physical parameters $\phi$, we express $\mathbf{K}_{\text{cont}}(\phi)$ as the sum of a static base dynamics and a parametric adjustment:
@@
Here, the base dynamics are governed by a learnable matrix $\mathbf{K}_{0}$, which captures the global behavior shared across the systems. $\mathcal{N}_{\psi}$ is Low-Rank Adaptation (LoRA (Hu et al., 2022)) parameterized by $\psi$, and it predicts the regime-specific deviations induced by $\phi$ and outputs a correction matrix matching the shape of $\mathbf{K}_{0}$.
@@ 4.2 Training
The training procedure follows a recurrent rollout scheme. Given an input sequence context $\{x_{t_{i-1}},x_{t_{i}}\}$, the model predicts a future trajectory of length $N$, denoted as $\{\hat{x}_{t_{i+1}},\hat{x}_{t_{i+2}},\dots,\hat{x}_{t_{i+N}}\}$. An architecture overview is provided in Figure 1, detailing the introduction of initial latent states, the latent dynamics propagation and how each of the states can be independently mapped back to the original data space.
@@
The rollout prediction loss penalizes errors over future time steps produced by iterating the learned latent dynamics and decoding the resulting latent states,
@@
where $N$ is the rollout horizon and $\{w_{i+j}\}_{j=1}^{N}$ are normalized temporal weights. In this work, the weights are either uniform or follow a cosine schedule that emphasizes early prediction accuracy,
$w_{i+j}\propto\tfrac{1}{2}\left(1+\cos\left(\tfrac{\pi(i+j-1)}{N-1}\right)\right).$ (14)
@@
In addition to the standard reconstruction and prediction losses, we include physics-inspired penalties that improve temporal alignment and structural fidelity of long rollouts. These include: (i) a temporal Sobolev loss matching finite-difference time derivatives, (ii) a spatial gradient loss enforcing edge and structure consistency, and (iii) a spectral loss computed in the Fourier domain to penalize frequency and phase mismatches. All physics terms are aggregated into a single auxiliary loss $\mathcal{L}_{\mathrm{phys}}$. More details can be found in Appendix C.
@@
where $\alpha$ and $\beta$ control the relative contributions of the rollout prediction and latent consistency losses, respectively, and $\lambda_{\mathrm{phys}}$ weights the optional physics-informed regularization terms.

### v2

@@ Abstract
Learning surrogate models for time-dependent PDEs requires balancing expressivity, stability, and computational efficiency. While highly expressive generative models achieve strong short-term accuracy, they rely on autoregressive sampling procedures that are computationally expensive and prone to error accumulation over long horizons. We propose a continuous-time Koopman autoencoder in which latent dynamics are governed by a parameter-conditioned linear generator. This formulation enables exact latent evolution via matrix exponentiation, allowing predictions at arbitrary temporal resolutions without autoregressive rollouts. We evaluate our method on challenging fluid dynamics benchmarks and compare against autoregressive neural operators and diffusion-based models. We evaluate our method on challenging fluid dynamics benchmarks against autoregressive neural operators and diffusion-based models. Our results demonstrate that imposing a continuous-time linear structure in the latent space yields a highly favorable trade-off: it achieves massive computational efficiency and extreme long-horizon stability while remaining competitive in short-term generative accuracy.
@@ 1 Introduction
Deep learning has recently emerged as a promising tool for flow prediction (Borrelli et al., 2022; Srinivasan et al., 2019; Fukami et al., 2020; Morimoto et al., 2021). Classical data-driven solvers operate directly on the data space, mapping flow fields at the current timestep to the next autoregressively using models such as Convolutional Neural Networks (CNNs) (Krizhevsky et al., 2012; Bhatnagar et al., 2019; Guo et al., 2016), Graph Neural Networks (GNNs) (Kipf and Welling, 2017; Rubanova et al., 2021; Pfaff et al., 2020) and Neural Operators (Lu et al., 2021; Li et al., 2021). While such data-space models excel at capturing fine-scale features, they tend to suffer from error accumulation (Brandstetter et al., 2022), high computational cost due to step-by-step rollouts, and reduced stability over long horizons.
Latent-space models compress high-dimensional flow fields into low-dimensional representations and evolve them over time using learned dynamics. Prior work (Vlachas et al., 2018; Wu et al., 2022) demonstrates that learning global latent dynamics can provide a simple and scalable alternative. A principled formulation of latent dynamics is given by Koopman theory (Koopman, 1931), which represents nonlinear systems through linear evolution in a lifted feature space. In practice, this operator is infinite-dimensional and must be approximated. Koopman autoencoders address this by learning a nonlinear mapping to a finite-dimensional latent space in which the dynamics evolve linearly (Takeishi et al., 2017; Lusch et al., 2018).
However, existing Koopman-based approaches typically learn discrete-time operators that advance the latent state by a fixed timestep, tying the learned dynamics to the temporal resolution of the training data. Inspired by continuous-time modeling approaches such as neural ODEs (Chen et al., 2018), we instead represent latent evolution using a linear Koopman generator. Unlike general neural ODE models, this formulation preserves a linear latent structure, enabling closed-form solutions through matrix exponentiation, interpretable spectral properties of the dynamics, and efficient long-horizon forecasting without iterative integration. In addition, we model this continuous operator as a function of governing physical parameters (e.g., Reynolds and Mach numbers), which enables a single model to capture families of dynamical systems across different regimes rather than learning a fixed operator tied to a single configuration.
Our approach focuses on deterministic latent dynamics designed to prioritize long-horizon stability, temporal flexibility, and computational efficiency. To rigorously evaluate these capabilities, we benchmark our model against a diverse set of baselines, including expressive generative models and data-space autoregressive predictors. Our evaluation demonstrates that enforcing a structured, continuous-time linear operator in the latent space yields significant advantages over highly expressive but unstable autoregressive methods. Ultimately, these findings establish continuous-time Koopman models as a highly scalable and robust alternative for long-horizon PDE forecasting.
@@
  • Continuous-Time Koopman Autoencoder: We formulate latent evolution as a continuous-time linear dynamical system governed by a learned Koopman operator, enabling closed-form integration via matrix exponential and eliminating the need for autoregressive rollouts.
  • Physics-Informed, Parameter-Conditioned Latent Dynamics: The Koopman operator is conditioned on physical parameters, allowing a single model to represent families of flow regimes while maintaining physically consistent predictions.
  • Temporal Resolution Super-Resolution: By modeling the latent dynamics in continuous time, the learned system can be evaluated at arbitrary timesteps, enabling zero-shot prediction at unseen temporal resolutions.
  • Efficient Long-Horizon Forecasting: The proposed formulation enables non-autoregressive prediction through matrix exponential evolution of the latent state, resulting in substantially faster inference compared to diffusion-based baselines.
@@ 2 Related Work
Autoregressive prediction in data space. Classical deep learning approaches for flow forecasting adopt autoregressive prediction in the data space. CNNs, particularly U-Net architectures (Ronneberger et al., 2015), have been widely used to capture complex spatial features in a dense image modeling framework (Deo et al., 2023; Wang et al., 2019). Other approaches have also explored generative models. For instance, Generative Adversarial Networks (GANs) (Goodfellow et al., 2020) have been employed to predict complex unsteady flow dynamics around objects like cylinders (Lee and You, 2019). Similarly, conditional diffusion models (Batzolis et al., 2021) have been benchmarked for their ability to predict turbulent flow simulations autoregressively (Kohl et al., 2026).
Latent-space autoregressive models. Many prior works compress flow fields into latent representations and evolve them using autoregressive dynamics. While this decouples spatial and temporal complexity, these approaches rely on step-to-step prediction, leading to error accumulation and limited temporal flexibility. Classical examples include CNN autoencoders coupled with LSTMs (Hochreiter and Schmidhuber, 1997) and Recurrent Neural Networks (RNNs) to propagate the latent state forward in time (Eivazi et al., 2020; Hasegawa et al., 2020; Bukka et al., 2020; Maulik et al., 2021; Zhou and Cheng, 2025). Similarly, Graph Neural Networks (GNNs) have proven effective for systems with irregular geometries to conduct temporal evolution in latent space (Sanchez-Gonzalez et al., 2020). More recently, transformers (Vaswani et al., 2017) have also been adapted for this task, leveraging its powerful attention mechanism to capture long-range temporal dependencies in the latent dynamics of fluid flows (Hemmasian and Farimani, 2023).
Koopman autoencoders. Building upon Lusch et al. (2018)’s foundational work, subsequent studies have introduced various enhancements. Azencot et al. (2020) proposed Consistent Koopman Autoencoders, integrating both forward and backward dynamics to enhance accuracy for long-time predictions. The Temporally-Consistent Koopman Autoencoder (tcKAE) (Nayak et al., 2025) achieved accurate long-term predictions with limited and noisy training data through a consistency regularization term. Halder et al. (2026) proposed Koopman $\beta$-variational autoencoders for reduced-order modelling (ROM) of turbulent flows, while $K^{2}$VAE integrates KalmanNet with KAEs to refine predictions and model uncertainty (Wu et al., 2025). Similar to our work, Menier et al. (2025) modeled continuous-time latent dynamics using a linear Koopman operator augmented by a nonlinear Mori-Zwanzig closure.
Different from existing works, we formulate the latent dynamics as a strictly continuous-time linear ODE. Crucially, we parameterize the Koopman operator as a function of governing physical inputs. This decoupled, continuous formulation allows for highly efficient integration at arbitrary temporal resolutions via matrix exponentiation, while the physical conditioning enables a single model to adapt and generalize seamlessly across different flow regimes.
@@ 3.1 Koopman Autoencoders
Koopman operator theory provides a linear perspective on nonlinear dynamics. Consider a discrete-time dynamical system with states $x_{t}\in\mathbb{R}^{N_{d}}$, where $x_{t}$ denotes the state of the system at time $t\in N^{d}$. The state evolves to its next timestep $t+\Delta t$ according to a flow map $f:\mathbb{R}^{N_{d}}\to\mathbb{R}^{N_{d}}$:
@@
which serves as a finite-dimensional approximation of the observable space. In this latent space, the evolution can be modelled as an approximately linear finite-dimensional Koopman operator $\mathbf{K}$,
@@
enabling evaluation at arbitrary timesteps and extrapolation beyond the training discretization.
@@ 4.1 Proposed Model
Let $x_{t}\in\mathbb{R}^{N_{d}}$ denote the high-dimensional physical state of the system. In our training datasets, the flow snapshots are provided at discrete timestamps $t_{i}$, uniformly spaced by a fixed time interval $\Delta t$, such that $t_{i+1}=t_{i}+\Delta t$. Although the formulation naturally supports irregularly sampled data, we use uniformly spaced timesteps in our experiments to align with standard benchmark protocols and enable direct comparison with existing baselines.
@@ 4.1.1 Encoder and Decoder
We employ a Transformer-based encoding architecture consisting of two symmetric streams: a history encoder ($\mathcal{E}_{\text{history}}$) and a present encoder ($\mathcal{E}_{\text{present}}$). To maintain consistency with benchmark evaluations, we formulate the temporal prediction as an Auto-Regressive process of order 2 (AR-2). Mathematically, we model the conditional distribution of the future state as $p(x_{t_{i+1}}|x_{t_{i}},x_{t_{i-1}})$.
Crucially, both encoders share identical weights and architecture, effectively acting as a single Siamese network that processes the immediate past state $x_{t_{i-1}}$ and the current state $x_{t_{i}}$ in parallel. The initial latent state $z_{t_{i}}$ is then constructed by averaging the feature representations from both streams:
@@ 4.1.2 Koopman Operator
Next, we discuss the structure of the Koopman operator. While conditioning neural surrogate models on physical parameters is common in PDE forecasting, incorporating this conditioning directly into a Koopman operator has received comparatively little attention. Therefore, a key innovation in our approach is explicitly conditioning the continuous-time Koopman operator on these external control parameters, denoted by $\phi$. In our setting, $\phi$ represents physical parameters such as Reynolds or Mach numbers. Unlike traditional Koopman approaches where a single static operator is learned for the entire domain, our operator adapts dynamically based on the specific physical regime.
@@
where fourth-order Runge–Kutta (RK4) is used to conduct this integration stably.
Finally, we define the structure of the operator. To capture the dependency on physical parameters $\phi$, we express $\mathbf{K}_{\text{cont}}(\phi)$ as the sum of a static base dynamics matrix and a parametric adjustment:
@@
Here, the base dynamics are governed by a learnable matrix $\mathbf{K}_{0}$, which captures the global, invariant behavior shared across all flow regimes. $\mathcal{N}_{\psi}$ is a Low-Rank Adaptation (LoRA (Hu et al., 2022)) module parameterized by $\psi$. We use a low-rank parameterization to control the operator spectrum and maintain numerical stability. This restricts $\mathbf{K}_{\text{cont}}$ from deviating too radically from the stable base dynamics $\mathbf{K}_{0}$, while significantly reducing the computational overhead of parameterizing an entire matrix for every distinct flow regime.
During training, we inject small Gaussian noise into the conditioning parameters $\phi$. This acts as a regularizer that improves robustness to sparse parameter sampling and discretization artifacts. In practice, this encourages the learned Koopman operator to vary smoothly across parameter space, improving interpolation between regimes.
@@ 4.2 Training
The training procedure follows a recurrent rollout scheme. Given an input sequence context $\{x_{t_{i-1}},x_{t_{i}}\}$, the model predicts a future trajectory of length $N$, denoted as $\{\hat{x}_{t_{i+1}},\hat{x}_{t_{i+2}},\dots,\hat{x}_{t_{i+N}}\}$. An architecture overview is provided in Figure 1.
@@
The rollout prediction loss penalizes errors over future time steps produced by iterating the learned latent dynamics and decoding the resulting latent states. This term mitigates error accumulation by training on multi-step rollouts.
@@
where $N$ is the rollout horizon and $\{w_{i+j}\}_{j=1}^{N}$ are normalized temporal weights. In this work, the weights follow a cosine schedule that emphasizes early prediction accuracy: $w_{i+j}\propto\frac{1}{2}\left(1+\cos\left(\frac{\pi(i+j-1)}{N-1}\right)\right)$. We use a decaying temporal weighting because autoregressive models are highly susceptible to compounding errors; enforcing stricter accuracy on immediate short-term predictions ensures the model establishes a correct base trajectory.
@@
In addition to the standard reconstruction and prediction losses, we include additional losses that enforce temporal and spatial consistency of long rollouts. These include: (i) a temporal Sobolev loss matching finite-difference time derivatives, (ii) a spatial gradient loss enforcing edge and structure consistency, and (iii) a spectral loss computed in the Fourier domain to penalize frequency and phase mismatches. All physics terms are aggregated into a single auxiliary loss $\mathcal{L}_{\mathrm{phys}}$. More details can be found in Appendix C.
@@
where $\alpha$ and $\beta$ control the relative contributions of the rollout prediction and latent consistency losses, respectively, and $\lambda_{\mathrm{phys}}$ weights the optional physics-informed regularization terms. More details about how the hyperparameters were chosen can be found in C.
@@
4.3 Inference and Integration via Matrix Exponentiation
A major advantage of restricting the latent dynamics to a linear Koopman form is that, at inference time, we can leverage the analytical solution of the learned system through matrix exponentiation. This bypasses iterative ODE solvers entirely and yields efficient long-horizon forecasting. Given the linear system in (8), the exact solution for any future state $z_{\tau}$ given $z_{0}$ is defined by the matrix exponential $z_{\tau}=\exp(\mathbf{K}_{\text{cont}}\tau)z_{0}$. This formulation allows us to predict the state at an arbitrary future time $\tau$ in a single computational step, independent of the training step size $\Delta t$. The matrix exponential $\exp(\mathbf{M})$ is defined by the power series: